OpenStudy (madhatter):

plz help The temperature on a particular day started at –13°F. It rose steadily by 3° each hour. The function y = –13 + 3x models the temperature, where x is the number of hours and y is the temperature. Graph the equation.

jimthompson5910 (jim_thompson5910):

y = –13 + 3x is the same as y = 3x - 13

jimthompson5910 (jim_thompson5910):

the second equation is in the form y = mx + b where m = 3 (slope) b = -13 (y-intercept)

jimthompson5910 (jim_thompson5910):

the y-intercept is -13, which means we know the point (0, -13) is on the graph

jimthompson5910 (jim_thompson5910):

the slope is 3 or 3/1 so this tells us to go up 3 units then go to the right one unit to land at the point (1, -10)

jimthompson5910 (jim_thompson5910):

So we have these two points now (0, -13) and (1, -10) Plot them and draw a straight line through them

OpenStudy (madhatter):

omg thanks so much

jimthompson5910 (jim_thompson5910):

you're welcome